### v3

@@ 5 Discussion
This study systematically investigates human mobility patterns in the Tokyo metropolitan area using a proposed analytical framework based on potential landscape and PCA. By comparing behaviors across distinct contexts, we identify clear distinctions in movement dynamics. Our method effectively captures differences in daily mobility patterns between weekdays and holidays, and reveals significant changes in mobility behavior during the COVID-19 pandemic. These findings provide a comprehensive understanding of how human mobility adapts to external factors, such as public health crises, across both temporal and spatial dimensions. Notably, the proposed approach leverages the potential landscape framework to represent and analyze the dynamics of human flow. This method serves as a powerful tool for visualizing mobility patterns by capturing the temporal evolution of spatial potential distributions. Mapping these shifts onto a dynamic landscape reveals intricate variations in behavior, such as constrained mobility during peak hours in 2021 and notable differences in holiday and weekday movements. Unlike static mobility data, the proposed approach offers deeper insight into the response of human movements to routine activities and external disruptions.
However, this study has several limitations. First, we relied on aggregated two-month mobility averages for 2019 and 2021, which restricted our ability to detect transient fluctuations, such as those driven by weather, short-term events, or sudden policy shifts. Second, we adjusted the GPS data using population census information, assuming proportional relationships between the sampled device users and the broader population. This assumption may overlook potential variations in local population density and disparities in device usage, thereby introducing bias. Furthermore, although the proposed method reproduced broad flow tendencies between regions, specific OD pair information was not retained. Such granularity can be critical for transportation planning, where absolute flow volumes between specific OD pairs are often essential.
@@
Nonetheless, the proposed method provides a macroscopic view of spatiotemporal mobility patterns, making it well-suited for applications that do not require a high level of detail. Our findings have direct relevance for various applications. A deeper understanding of mobility patterns can inform urban planning, support the design of optimized transportation networks, and guide the allocation of public resources. From a public health perspective, insights from potential landscapes can help policymakers develop targeted interventions during pandemics and other crises. Additionally, the potential landscape approach provides policymakers a dynamic tool for tracking mobility trends, enabling them to adjust infrastructure and regulations in response to evolving urban needs. Future studies with higher temporal resolution and real-time data are essential for expanding the scope of these applications and enhancing the precision of mobility analyses.
@@
The authors declare that they have no competing interests.
@@
The mobility flow data for Japan utilized in this study is available for purchase from LocationMind Inc., a Japanese company [68] (For inquiries, use their contact form [72]). The data product is called “LocationMind xPop.”
@@ Acknowledgements
Y.D. is supported by JST SPRING, Grant Number JPMJSP2114. T.A. is supported by JSPS KAKENHI Grant Number JP24K07699, and JSPS KAKENHI Grant Number JP24H01473. N.F. is supported by JSPS KAKENHI Grant Number JP24K03007, and JST PRESTO Grant Number JPMJPR21RA, Japan. This work additionally receives support from the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
Supplementary Information
@@
To investigate the dependency of mobility flow on distances, Figure 6 shows the percentage of non-zero mobility pairs ($M_{ij}>0$ or $M_{ji}>0$) as a function of road distance in the Japanese mobility dataset [68]. A rapid decrease is observed as the distance increases. Notably, even at a relatively short distance of approximately $d_{ij}\sim 40$ km, around half of the mobility pairs exhibit zero flow volume.
Figure 6: Percentage of OD pairs $(i,j)$ with non-zero mobility as a function of spatial distance $d_{ij}$. In all range of distance, number (left) and percentage (right) of non-zero mobility pairs are plotted.
Supplementary Note 2: A threshold weighting method
We employed a threshold weighting method to determine the edges $E$ following [56], as discussed in the main text. Specifically,the threshold $\theta$ was established as the distance below which the majority of trips occur, based on the probabilistic criterion $p(X<\theta)=0.99$. Only the edges with distances below the threshold were considered in this work. The threshold distances for the four different scenarios are shown in Fig. 7.
Figure 7: Threshold distances $\theta$ for the four scenarios: (a) 2019 Weekday; (b) 2010 Holiday; (c) 2021 Weekday; and (d) 2021 Holiday.
Supplementary Note 3: Variances explained by principle components
We conducted a PCA on the combined data from weekdays and holidays in both 2019 and 2021. Notably, the first three principle components capture the majority of the variance, as shown in Fig. 8. This indicates that they play a crucial role in explaining the underlying mobility patterns.
Figure 8: Screeplot of the PCA results for the combined data from weekdays and holidays in 2019 and 2021. The first three principle components capture the majority of the variance.
@@
The hourly detailed potential landscapes for the 2019 weekday scenario are shown in Figs. 9 and 10. The potential landscapes exhibit a clear pattern as discussed in the main text.
(a) $0$ 01:00
(b) $0$ 01:00
@@
The background map layouts in Figs. 9 and 10 are based on data by © OpenStreetMap contributors, available under the Open Database License (ODbL). Map tiles by © CARTO. The results in Figs. 6, 7, 8, 9, and 10 were obtained by using data “LocationMind xPop © LocationMind Inc” [68].

### v4

@@ 5 Discussion
This study systematically analyzes human mobility patterns in the Tokyo metropolitan area using a proposed analytical framework based on potential landscape and PCA. By comparing behaviors across distinct contexts, we identify salient differences in movement dynamics. Our method effectively captures differences in daily mobility patterns between weekdays and holidays and detects substantial shifts in mobility behavior during the COVID-19 pandemic. These findings elucidate how human mobility adapts to external factors, such as public health crises, across both temporal and spatial dimensions. Notably, the proposed approach uses the potential landscape framework to represent and analyze the dynamics of human flow. This method serves as an effective tool by capturing the temporal evolution of spatial potential distributions. Mapping these shifts onto a dynamic landscape reveals fine-grained variations in behavior, such as constrained mobility during peak hours in 2021 and notable differences in holiday and weekday movements. Unlike static mobility data, the proposed approach offers deeper insight into the response of human movements to routine activities and external disruptions.
Nevertheless, this study has several limitations. First, while the proposed reduction method is suitable for generic timestamped origin–destination datasets, we have only studied the Tokyo dataset before and during COVID-19. Applications to other cities or other contexts remain for future work. Second, notably, the potential landscape derived by the combinatorial Hodge theory does not capture the information of cycle flows between locations. As illustrated in Fig. 2, this reduction extracts the gradient component of a given human flow while omitting the other cyclic components. Unlike the gradient component, such cyclic components do not induce temporal changes in the local population, because the incoming and outgoing fluxes are balanced at each place. In other words, the proposed method only focuses on the imbalanced population flows toward higher-potential places from lower-potential places. Third, the proposed method cannot directly identify trip purpose or activity types from the data. For example, while we infer that specific patterns represent commuting, this is an inference based on spatiotemporal evidence. This issue warrants future studies. One approach is to leverage person-trip survey data with travel purpose or activity types. Another solution is integrating GPS-tracked mobility data with external point-of-interest datasets to infer the activity more directly.
@@
Nonetheless, the proposed method offers a macroscopic view of spatiotemporal mobility patterns, rendering it well-suited for applications that do not require a high level of detail. Our findings have direct relevance for several applications. A deeper understanding of mobility patterns can inform urban planning, support the design of efficient transportation networks, and guide the allocation of public resources. From a public health perspective, insights from potential landscapes can help policymakers develop targeted interventions during pandemics and other crises. Additionally, the potential landscape approach offers decision-makers a dynamic tool for tracking mobility trends, enabling them to adjust infrastructure and regulations in response to evolving urban needs. Future studies with higher temporal resolution and real-time data will be crucial for expanding the scope of these applications and enhancing the precision of mobility analyses.
@@
The authors report no conflicts of interest.
@@
The mobility flow data for Japan utilized in this study are available for purchase from LocationMind Inc., a Japanese company [LocationMind] (for inquiries, use their contact form [LocationMind_contact]). The data product is named “LocationMind xPop.”
@@ Acknowledgements
Y.D. is supported by JST SPRING, Grant Number JPMJSP2114. T.A. is supported by JSPS KAKENHI Grant Number JP24K07699, and JSPS KAKENHI Grant Number JP24H01473. N.F. is supported by JSPS KAKENHI Grant Number JP24K03007, and JST PRESTO Grant Number JPMJPR21RA, Japan. This work also receives support from the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
Supplementary Notes
@@
To investigate the dependence of mobility flow on distance, Figure 9 shows the percentage of non-zero mobility pairs ($M_{ij}>0$ or $M_{ji}>0$) as a function of road distance in the Japanese mobility dataset [LocationMind]. A rapid decline is observed as the distance increases. Notably, even at a relatively short distance of approximately $d_{ij}\sim 40$ km, around half of the mobility pairs exhibit zero flow volume.
Figure 9: Percentage of OD pairs $(i,j)$ with non-zero mobility as a function of spatial distance $d_{ij}$. In the range of distance, number (left) and percentage (right) of non-zero mobility pairs are plotted.
Supplementary Note 2: Sensitivity analysis by testing different values of $\theta$
In the first reduction step, we applied a weighting method with a threshold distance $\theta$ [AFF2023]:
$W_{ij}=\begin{cases}1\quad&\text{mobility pair $\{(i,j)\mid d_{ij}\leq\theta\}$}\\ 0\quad&\text{mobility pair $\{(i,j)\mid d_{ij}>\theta\}$}\end{cases}.$
The threshold $\theta$ is the distance below which the majority of trips occur, based on the probabilistic criterion $p(X<\theta)=0.99$. This criterion was chosen to include the majority of trips, while excluding possible outliers.
Table 2 shows a sensitivity analysis by testing different criteria ($p$ = 0.95 and 0.9), which led to different thresholds $\theta$. These results demonstrate that the resulting potential landscapes with modified parameters were very similar, maintaining high correlations even when the threshold $\theta$ was changed significantly.
Table 2: Correlation between the generated potential landscapes with modified pp.
Year Condition $p$ $p$ (modified) $\theta$ $\theta$ (modified) $\theta$ (changed) Correlaton
2019 Weekday 0.99 0.95 32.54 19.79 -39% 0.92
2019 Holiday 0.99 0.95 22.58 12.62 -44% 0.83
2021 Weekday 0.99 0.95 30.50 18.05 -41% 0.92
2021 Holiday 0.99 0.95 19.36 11.53 -40% 0.90
2019 Weekday 0.99 0.9 32.54 14.31 -56% 0.87
2019 Holiday 0.99 0.9 22.58 9.09 -60% 0.89
2021 Weekday 0.99 0.9 30.50 13.08 -57% 0.87
2021 Holiday 0.99 0.9 19.36 8.68 -55% 0.76
Supplementary Note 3: Variances explained by principal components
We conducted a PCA on the combined data from weekdays and holidays in both 2019 and 2021. Notably, the first three principal components capture the majority of the variance, as shown in Fig. 10. This indicates that they play a crucial role in explaining the underlying mobility patterns.
Figure 10: Screeplot of the PCA results for the combined data from weekdays and holidays in 2019 and 2021. The first three principal components capture the majority of the variance.
@@
The hourly detailed potential landscapes for the 2019 weekday scenario are shown in Figs. 11 and 12. The potential landscapes exhibit a clear pattern as discussed in the main text.
(a) 00:00 - 01:00
(b) 00:00 - 01:00
@@
The background map layouts in Figs. 11 and 12 are based on data by © OpenStreetMap contributors, available under the Open Database License (ODbL). Map tiles by © CARTO. The results in Figs. 9, 10, 11, and 12 were obtained by using data from “LocationMind xPop © LocationMind Inc” [LocationMind].